OpenStudy (anonymous):

Find the derivative of f(x) = -7/x at x = -3.

OpenStudy (anonymous):

I know Liebniz Notation I just can't figure out how to turn the -7/x into x^n

OpenStudy (bibby):

\(\large \frac{d}{dx}\frac{-7}{x}=-7\frac{d}{dx}\frac{1}{x}\)

OpenStudy (bibby):

than you can rewrite it as \(\large x^{-1}\)

OpenStudy (bibby):

then*

OpenStudy (anonymous):

that doesn't make any sense, sorry... Can you please try explaining it differently?

OpenStudy (bibby):

you can either use the quotient rule, or factor out the -7. There's a rule that states that \(\huge \frac{d}{dx}Cf(x)=C*\frac{d}{dx}f(x)\)

OpenStudy (bibby):

basically \(\large \frac{-7}{x}=-7*\frac{1}{x}\)

OpenStudy (anonymous):

so the answer is 7/3?

OpenStudy (bibby):

well first we need to find f'(x)

OpenStudy (anonymous):

Here's what I did: f(x)=-7/x at x=-3 f(x)=-7(1/x) at x=-3 f(x)=-7(1/-3) f(x)=2(1/3) f(x)=7/3 because the multiple choice answers are all improper fractions

OpenStudy (bibby):

The question asked for the derivative at x=-3 first we find the derivative next we plug in -3

OpenStudy (anonymous):

the derivative would be -7(1/x) at x=-3 or -7(1/-3)

OpenStudy (bibby):

no. the derivative would be \(\large -7\frac{d}{dx}\frac{1}{x}\)

OpenStudy (anonymous):

what's d?

OpenStudy (bibby):

d/dx is (leibniz?) notation for derivative

OpenStudy (anonymous):

ohh I learned Liebniz Notation as f(x)=x^n becomes nx^(n-1)

OpenStudy (bibby):

that's not what leibniz notation is. that's the power rule. f'(x) is another way of saying (first) derivative f(x)||f'(x) \(x^n\)||\(nx^{n-1}\)

OpenStudy (bibby):

so here we have something of the form \(x^{-1}\)

OpenStudy (anonymous):

okay, then what?

OpenStudy (bibby):

apply the power rule...

OpenStudy (anonymous):

-1(1)^(-1-1)=-1^-2=-1

OpenStudy (bibby):

why is there a 1 in parentheses? "-1(1)"

OpenStudy (anonymous):

to show the step

OpenStudy (bibby):

you're putting in numbers instead of letters and that makes no sense.

OpenStudy (bibby):

x^n -> nx^(n-1) \(\large x^{-1}=>-1*x^{-1-1}=-x^{-2}\)

OpenStudy (anonymous):

using the power rule, we start with x^n. You said we have x^-1. That means x=1 and n=-1. The second half of the power rule is nx^(n-1). Plug in n and x to get -1(1)^(-1-1) or -1^-2 which equals -1

OpenStudy (bibby):

that doesn't mean x=1. it means x=x

OpenStudy (bibby):

You need to brush up on your derivative table

OpenStudy (anonymous):

I just learned this...

OpenStudy (anonymous):

so if it's -x^-2, then what do I do?

OpenStudy (bibby):

I'm not trying to be harsh, all I'm saying is you'r given a table and a recipe, and you're making this stuff up instead of following the recipe.

OpenStudy (bibby):

you rewrite it in full. \(\huge -7*\frac{-1}{x^2}\) \(\huge f'(x) = \frac{d}{dx}=\frac{7}{x^2}\)

OpenStudy (anonymous):

so now we plug in x?

OpenStudy (bibby):

into the derivative equation, yeah

OpenStudy (anonymous):

so it's 7/9?

OpenStudy (bibby):

yeah

OpenStudy (anonymous):

can you maybe try to help me with one more?

OpenStudy (bibby):

post another question. we'll see